OpenStudy (anonymous):

how about this Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $25.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? 100 1000 500 50

OpenStudy (anonymous):

\[25 + 0.06x = 30 + 0.05x\] where x is the number of minutes.

OpenStudy (anonymous):

do you have to solve like terms?

OpenStudy (anonymous):

Answer: 50 .6x+25=.5x+30 -.5x -.5x ------------------------ .1x+25=30 -25 -25 ------------------ .1x=5 x=50 OR $0.05 x 500 = $25 + $30 = $55 $0.06 x 500 = $30 + $25 = $55 500 minutes. help

OpenStudy (anonymous):

What one is it

OpenStudy (anonymous):

yes...solve like like terms!

OpenStudy (anonymous):

its 500 because its 0.06 & 0.05 not 0.6 & 0.5